A manufacturer trying to get a process under control randomly collects 4 samples from the process each hour. Five of the X bar and R measurements are as follows: X bar: 374, 376, 360, 359, 369 R: 3, 2, 5, 2, 3 What are the control limit values for the two control charts (3 per chart)?
Answer #1

n : Number of observations in each group = 4

k: number of groups = 5

Given

Group \overline{X} R
1 374 3
2 376 2
3 360 5
4 359 2
5 369 3
Total \Sigma \overline{X}=1838 \Sigma R=15
\overline{\overline{X}} =\Sigma \overline{X}/k = 1838/5= 367.6 \overline{R}= \Sigma R /k =15/5 =3

Control limits for \overline{X} chart

Upper control Limit : UCL = X + AR

X = 10

Lower Control Limit : UCL = X - AR

A2 for n:4 = 0.729

UCL = X + AR= 367.6 + 0.729 X 3 = 367.6 +2.187 = 369.787

9298 = X = 10

LCL = X - AR = 367.6 - 0.729 X 3 = 367.6 – 2.187 = 365.413

Upper control Limit = 369.787

CL = 367.6

Lower Control Limit = 365.413

Control Limits R-Chart

UCL = DAR

= 15

LCL = D3R

For n:4 : D4 =2.282 ; D3 = 0

UCL = DAR= 2.282 x 3 = 6.846

CL = R=3

LCL = DR = 0 x 3 = 0